Circular No. 6198

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) R. H. McNaught, Anglo-Australian Observatory, reports that he has

located an image of this comet on a 50-min U.K. Schmidt R survey plate obtained by C. P. Cass on 1993 Apr. 27. During the exposure the comet, then 13.1 AU from the sun, would have moved 3′′.7 to the southwest, and the image is consistent with this. It has the appearance of a slightly diffuse and slightly elongated star image within a very faint coma 0′.4 in diameter. The total and nuclear magnitudes are ∼ 18 and 19, respectively, based on this being a good-quality plate with a presumed limit of 21. The position is 9′ from that given by the parabolic elements on MPEC 1995-P01 (and IAUC 6194), which are now improved on MPEC 1995-P02:

Epoch = 1995 Oct. 10.0 TT

T = 1997 Apr. 1.3922 TT  ω = 130.^{o}4405
e = 0.996348  Ω = 282.4733 }2000.0
q = 0.916702 AU  i = 88.8797

Epoch = 1997 Mar. 13.0 TT

T = 1997 Apr. 1.6416 TT  ω = 130.^{o}6678
e = 0.994441  Ω = 282.4729 }2000.0
q = 0.913023 AU  i = 89.4142

These orbital elements, satisfying 248 observations with mean residual 0′′.6, lead to (1/a)_{orig} = +0.00456 ± 0.00016 AU^{−1}, indicating that the comet is not on its first pass from the Oort Cloud. However, McNaught adds that he could not locate the comet on a plate taken on 1991 Sept. 1, when an inverse-square law would have put the comet 1.0 mag fainter than in 1993.

RINGS OF SATURN I. de Pater and J. R. Graham, University of California at Berkeley; J. J.

Lissauer, State University of New York at Stony Brook; and M. Showalter, NASA Ames Research Center, report: “Observations were made of Saturn’s rings on May 23 (1.3 days after the ring-plane crossing) with the W. M. Keck telescope on Mauna Kea at a wavelength of 2.3 μm. The E-ring is confined to within 2500 km from the ring plane; it extends out to 5 Saturnian radii. The peak brightness is ∼ 19 mag per linear arcsec at the orbit of Saturn II (Enceladus, 4 Saturnian radii). The ring brightness is lower and drops off faster with distance than seen at visible wavelengths. We also report a tentative detection of the G-ring at 2.8 Saturnian radii, similar in magnitude to the peak of the E-ring.”

1995 August 2 Brian G. Marsden